The inverse market demand in a homogeneous-product Cournot duopoly is P = 200 − 3(Q1 + Q2) and costs are C1(Q1) = 26Q1 and C2(Q2) = 32Q2. If the two oligopolists formed a cartel whose MC is the average between the oligopolists’ marginal costs, find the optimal output, price, and maximized profit. . Assume the cartel splits profit equally among oligopolists.
Answer #1

Average Costa 267 32 2 a S8= 294 2 7 = (200-34) Ų co - 254 da dů 200 3X20 29=0 Q = 28.5 P = 200-3X 28.5=114.5 7= (200 - 3X28.